OpenStudy (anonymous):

Find the molarity of 20.5 g of NaCl in 669 mL of solution.

OpenStudy (anonymous):

Molarity=no.of moles/1L thus no. of moles= 58.5/20.5=2.854 moles 1ml=1/1000L so 669ml = 0.669L Thus Molarity=2.854/0.669 = 4.27M